# Fermion-qubit fault-tolerant quantum computing

### v1

@@
By contrast, in our fermionic platform, the fact that fSWAP gates are implemented by atom rearrangements and can therefore be contracted (Fig. 4c), leads to a Clifford complexity of $\mathcal{O}(N\log(N))$ and a depth of $\mathcal{O}(\log N)$, yielding an exponential improvement for the latter. Moreover, our compilation of the constituent FFFT subroutines (see Fig. 4e, f) only takes two braiding and three single-fermion gates and T-gate depth 1, which is more gate-efficient than currently known qubit decompositions taking three CNOTs, nine single-qubit gates and T-gate depth 2 [13].
For the more general $H_{\mathrm{qchem}}$, at first sight, no depth advantage is possible because fermi-SWAP networks achieve the optimal depth $\mathcal{O}(N^{3})$ [14], which our approach reproduces with a smaller prefactor. However, asymptotically, there are only $\mathcal{O}(N^{2})$ non-negligible terms in the often-used Gaussian basis [41]. We therefore expect that, for large $N$, we will again yield at least an $\sim N$ depth advantage by the fact that our fermion approach always achieves the optimal depth by parallelization, whereas in the qubit approach, most of the fermi-SWAP network would still have to be traversed.
@@
Within our fault-tolerant gate set, no advantage in the number of T gates is possible if the T-gate count is the only parameter which is optimised for. This can be directly seen by mapping the fermionic gates to qubit gates using the Jordan-Wigner transform and noting that only the $\mathrm{T}^{\mathrm{f}}$ gate requires T gates (Methods). We note that steady progress in T-gate synthesis means they do not completely dominate the simulation cost over Clifford gates [30].
@@
Figure 5: Implementation of the braiding gate in neutral atoms. a, Illustration of the sequence of operations to implement an effective braiding gate. Braiding is a product of pairing and the tunneling gate $\sqrt{\mathrm{iSWAP}}^{\mathrm{f}}=U_{t\downarrow}(\pi/4)$. $U_{\mathrm{t}\downarrow}(\theta)$ and $U_{\uparrow\downarrow}(\theta)$ act only on fermions, while the beamsplitter acts only on molecules, and $U_{\mathrm{diss}}$ is a fermion-molecule gate. The pairing gate dissociates a bosonic molecule into two fermions on two separate sites. The pulse duration of the dissociation interaction depends on the average number $N=\braket{b^{\dagger}b}$ of molecules in the tweezer. b, The molecules are formed from the same atomic species as the fermionic atoms used. c, Ramsey sequence to probe the effective coherence of a “Bell-pair” (|$(\ket{00}+i\ket{11})/\sqrt{2}$ created by the effective pairing gate. The phase gate is defined as $\exp(-i\theta n)$. The beamsplitter gate, defined above Eq. (9), has phase $\theta=\pi/4$. d, Ramsey fringes as a function of initial molecule Fock state number $N$. e, Simulated pairing gate infidelity as a function of molecule Fock state number $N$. See Fig. S1f for the circuit used to determine the infidelity.
@@
Finally, we describe how to implement our framework in neutral atoms, which requires the implementation of the univeral qubit-fermion gateset shown in Fig. 1. $\mathrm{CZ}^{\mathrm{qf}}$ and $\mathrm{CZ}^{\mathrm{f}}$ are implemented using a high-lying Rydberg state of the atom, similarly to the routinely implemented qubit-qubit $\mathrm{CZ}^{\mathrm{q}}$ [42, 27]. $\sqrt{\mathrm{iSWAP}}^{\mathrm{f}}$ is implemented by bringing two tweezers close to each other, inducing tunneling of atoms between the sites [27] or by placing the atoms into an optical lattice [43] and controllably performing tunneling by using superlattices [44]. $\mathrm{T^{\mathrm{f}}}$, $\mathrm{S^{\mathrm{f}}}$, $\mathrm{Z^{\mathrm{f}}}$ are implemented by detuning the trapping frequency of the target tweezer or optical lattice site. What is left to show is how to implement $\mathrm{BRAID}$. We propose its realisation as a product $\mathrm{BRAID}_{ij}=\mathrm{PAIR^{\prime}}_{ij}\mathrm{\sqrt{iSWAP}}^{\mathrm{f}}_{ij}$, where $\mathrm{PAIR^{\prime}}_{ij}$ is the pairing gate
@@
To implement pairing, we propose the pulse-sequence in Fig. 5a. It employs photodissociation of homonuclear diatomic bosonic molecules consisting of the same atomic species as the fermions [45], see Fig. 5b. Photodissociation realizes the boson-fermion gate $U_{\mathrm{diss}}=\exp\left(-i\frac{\pi}{4\sqrt{N}}\left(b_{i}p^{\dagger}_{i\uparrow}p^{\dagger}_{i\downarrow}+b^{\dagger}_{i}p_{i\downarrow}p_{i\uparrow}\right)\right)$, where $b$ is the bosonic annihilation operator of molecules (we assume there are $N$ of them) residing in the ground state of the tweezer and obeying commutation relations $[b_{i},b_{j}^{\dagger}]=\delta_{ij}$. $\uparrow$ and $\downarrow$ label two internal states of the atom, where so far we assumed that all atoms reside in $\downarrow$, i.e. $p_{i}=p_{i\downarrow}$. In order to convert this same-site gate into a non-local gate, we conjugate it with a fermion spin-selective tunneling $U_{\mathrm{t}\downarrow}(\theta)=\exp\left(-i\theta\left(p^{\dagger}_{i\uparrow}p_{j\uparrow}+h.c.\right)\right)$ and a fermionic spin-flip $U_{\uparrow\downarrow}(\theta)=\exp\left(-i\theta\left(p^{\dagger}_{i\downarrow}p_{i\uparrow}+h.c.\right)\right)$. To suppress interactions between the molecules and the fermions, the molecules reside in a separate tweezer when not in use for gates and are tunneled into the target tweezer by moving the two tweezers close to each other, implementing a beam-splitter operation $\mathrm{BS}(\theta)=\exp\left(\theta\left(b_{i}^{\dagger}b_{j}-h.c.\right)\right)$. Altogether, the sequence in Fig. 5a implements
$\mathrm{PAIR}_{ij}=\exp\left(i\frac{\pi}{4\sqrt{N}}\left(b_{k}p_{i}^{\dagger}p_{j}^{\dagger}+h.c.\right)\right).$ (9)
@@
We now discuss how we use Eq. (9) to effectively implement Eq. (8). If the molecules were in a coherent state with large amplitude $\sqrt{N}$, we could replace $b\rightarrow\braket{b}=\sqrt{N}$ and the two gates would be equal. However, in the case of atoms trapped in dipole traps, Fock states or mixtures thereof are more naturally prepared than coherent states. To overcome this conceptual challenge, we use the little-known fact that large-$N$ Fock states cannot be distinguished from coherent states by measurements [46].
@@
To show that we can indeed effectively implement Eq. (8) employing a large-$N$ Fock state of molecules, consider the thought experiment shown in Fig. 5c. The first pairing gate is supposed to create a “fermionic Bell-pair” (|$(\ket{00}+i\ket{11})/\sqrt{2}$. To probe whether there is coherence, a phase gate with phase $\theta$ is applied and a second $\mathrm{PAIR}$ gate rotates the measurement basis into the effective Bell basis. This realizes a Ramsey-type experiment. Simulating this experiment numerically, we find Ramsey fringes with a contrast that approaches unity as $N\rightarrow\infty$. This shows that, after the first pairing gate, there was indeed effective coherence between states |$\ket{00}$ and |$\ket{11}$. An intuitive way to understand this observation is that the molecule number distribution in each mode after the beamsplitter is approximately Poissonian—just like in a coherent state. The entanglement between both modes then leads to an effective phase coherence between the modes. This effective coherence is then revealed by the second pulse, which disentangles the fermions from the molecules. We say “effective” coherence because the reduced density matrix of the fermions is always diagonal; indeed, by measuring the total number of molecules, one projects the fermions onto |$\ket{00}$ or |$\ket{11}$. Therefore, it is important that the molecules are not measured while the fermions are entangled with them. In Fig. S1, we additionally perform state tomography numerically, showing that the state after the first pairing gate is indeed effectively a Bell pair.
@@
We probe the gate fidelity of our effective pairing gate in Fig. 5e using Choi state tomography, which we generalize to fermionic gates, see circuit shown in Fig. S1. We find that only $100$ molecules suffice to create a gate with $>99.8\,\%$ fidelity. In the Methods, we also show that even a mixed state with a Poissonian number distributation can be used instead of |$\ket{N}$. Non-trivially, Fig. 5c and Fig. 5d show that our scheme works even if the second pairing gate is not performed on the same molecule site, which enables full parallelizability of pairing operations. For that to be the case, the molecule tweezers need to be effectively coherent with respect to each other, just like lasers manipulating different qubits within a quantum computer need to be phase-locked. The cross-coherence of molecule tweezers is guaranteed by creating them from one common Fock state with a beamsplitter as done in Fig. 5c. To parallelize this further, many molecule tweezers can be created by repeatedly applying beamsplitters starting from one large Fock state, for example by tunneling from a single large dipole trap.
@@
Our work opens the field of fermion-qubit error correction and fault-tolerant algorithms. Of prime importance is the error threshold of our codes for the strongly biased errors present in experiments. Moreover, fermionic versions of quantum low-density-parity check codes may be developed to enable a finite rate of logical fermions per codeblock. Due to our reliance on transversal gates, algorithmic fault tolerance can be explored within our approach [30]. Finally, we assumed that the bosonic molecule mode does not need to be error-corrected, which might require a more restrictive lower bound on the mode occupation $N$.
Our present scheme employs single-species fermions. While two-species models, for example spin up and spin down electrons in quantum chemistry, can be mapped to single-species computation, an architecture involving two species, as partially employed in our pairing gate, might make the implementation more efficient.
@@
Finally, our architecture might offer an advantageous approach not only for simulating fermions but also for quantum computations that are natively qubit-only. We encoded logical fermion operators in odd-weight fermion operators—instead, logical qubits can be encoded in even-weight logical operators, for example Majorana surface codes which have been shown to have high thresholds [59]. One advantage of this approach is that lost atoms are replenished by coupling to the molecule bath, offering an alternative to continuous atom reloading [43], which is one of the challenging aspects of neutral-atom quantum computing.
@@ Acknowledgments
We acknowledge discussions with Christopher Kang, Paul Julienne, Robert Ott, Shraddha Singh, Maryam Mudassar, Nathan Wiebe, Torsten Zache, Johannes Zeiher, and Hengyun Zhou.

### v2

@@
By contrast, in our fermionic platform, the fact that fSWAP gates are implemented by atom rearrangements (Fig. 4c) means that a depth of $\mathcal{O}(\log N)$ can be implemented without ancillas, yielding $\mathcal{O}(N)$ space and $\tilde{\mathcal{O}}(N)$ volume complexity. We note that the $N$-scaling of movement time does not prevent the exponential depth reduction from leading to an exponential time saving. This is because the constant overhead of error correction dominates, see Methods. We achieve another depth improvement through our compilation of the constituent FFFT subroutines (see Fig. 4e, f), which only takes half the previously-known T-gate depth [13].
At first sight, the logarithmic depth of the kinetic term is not useful because the interaction term requires $\mathcal{O}(N)$ depth because of the parallelizability restriction imposed by the $\mathcal{O}(N^{2})$ terms. However, similarly to the qubit version of the FFFT, we can use ancillas to reduce the circuit depth by using the translational invariance of $U_{ij}$. This enables using a quantum circuit of the classical fast Fourier transformation [46, 49], see Methods. Because the Fourier transform of $n_{i}$ is a complex number, we encode it in a binary register with $\mathcal{O}(\log(N))$ ancillas. Therefore, this “parallelized interactions” scheme in total uses $\mathcal{O}(N\log(N))$ ancillas, for both qubit and fermion quantum computers. After this transform, the density interactions are diagonal and implementable in depth $\mathcal{O}(\mathrm{log}(N))$.
Taking both interaction and kinetic terms together, we thus achieve a log-depth Trotter step for materials. For qubit quantum computers, it uses $\mathcal{O}(N^{2})$ qubits, but for fermionic quantum computers only $\mathcal{O}(N\log(N))$ fermionic sites. Therefore, the circuit volume is a factor of $N/\log(N)$ lower for fermions.
The remaining question is how this log-depth Trotter step reflects on the end-to-end complexity of time-evolving up to time $t$. We use the Trotter scheme introduced in Ref. [50], which uses $\mathcal{O}(Nt)$ Trotter steps for fixed simulation accuracy and large, fixed Trotter order. Using our log-depth Trotter steps, we therefore find an overall depth complexity of $\tilde{\mathcal{O}}(Nt)$, which is essentially linear in both time and the number of sites. While we show that both qubit and fermion quantum computers can reach this scaling, only the fermion version requires a linear number of modes, while at least within our approach, qubits need a prohibitively expensive quadratic number of qubits using fSWAP networks.
For the more general $H_{\mathrm{qchem}}$, at first glance, no depth advantage appears possible, since current state-of-the-art schemes combining double factorization in the molecular basis with fSWAP networks achieve depth $\mathcal{O}(N^{2})$ without ancillas [51] and $\mathcal{O}(N)$ depth with ancillas [52] for a single Trotter step, independent of the molecule size. By contrast, fermionic quantum computers can reach depth $\mathcal{O}(N)$ for large molecules, see Methods, which to our knowledge is the first algorithm for time evolution under the chemistry Hamiltonian which achieves linear depth per Trotter step without using ancillas. For Hamiltonians which do not allow for a low-rank factorization, such as the sparse Sachdev-Ye-Kitaev model [53], we expect such depth speedups to become apparent already at small system sizes.
@@
Within our fault-tolerant gate set, no advantage in the number of T gates is possible if the T-gate count is the only parameter which is optimised for. This can be directly seen by mapping the fermionic gates to qubit gates using the Jordan-Wigner transform and noting that only the $\mathrm{T}^{\mathrm{f}}$ gate requires T gates (Methods). We note that steady progress in T-gate synthesis means they do not completely dominate the simulation cost over Clifford gates [34]. Moreover, our depth reductions directly translate into a large space-time volume saving compared to qubit platforms. Therefore, we expect advantages in terms of the total runtime of a fermionic simulation.
@@
Figure 5: Implementation of the braiding gate in neutral atoms. a, Illustration of the sequence of operations to implement an effective braiding gate. Braiding is a product of pairing and the tunneling gate $\sqrt{\mathrm{iSWAP}}^{\mathrm{f}}=U_{t\downarrow}(\pi/4)$. $U_{\mathrm{t}\downarrow}(\theta)$ and $U_{\uparrow\downarrow}(\theta)$ act only on fermions, while the beamsplitter acts only on molecules, and $U_{\mathrm{diss}}$ is a fermion-molecule gate. The pairing gate dissociates a bosonic molecule into two fermions on two separate sites. The pulse duration of the dissociation interaction depends on the average number $M=\braket{b^{\dagger}b}$ of molecules in the tweezer. b, The molecules are formed from the same atomic species as the fermionic atoms used. c, Ramsey sequence to probe the effective coherence of a “Bell-pair” (|$(\ket{00}+i\ket{11})/\sqrt{2}$ created by the effective pairing gate. The phase gate is defined as $\exp(-i\theta n)$. The beamsplitter gate, defined above Eq. (9), has phase $\theta=\pi/4$. d, Ramsey fringes as a function of initial molecule Fock state number $M$. e, Simulated pairing gate infidelity as a function of molecule Fock state number $M$. See Fig. S1f for the circuit used to determine the infidelity.
@@
Finally, we describe how to implement our framework in neutral atoms, which requires the implementation of the universal qubit-fermion gateset shown in Fig. 1. $\mathrm{CZ}^{\mathrm{qf}}$ and $\mathrm{CZ}^{\mathrm{f}}$ are implemented using a high-lying Rydberg state of the atom, similarly to the routinely implemented qubit-qubit $\mathrm{CZ}^{\mathrm{q}}$ [54, 36]. $\sqrt{\mathrm{iSWAP}}^{\mathrm{f}}$ is implemented by bringing two tweezers close to each other, inducing tunneling of atoms between the sites [36] or by placing the atoms into an optical lattice [55] and controllably performing tunneling by using superlattices [56, 57]. $\mathrm{T^{\mathrm{f}}}$, $\mathrm{S^{\mathrm{f}}}$, $\mathrm{Z^{\mathrm{f}}}$ are implemented by detuning the trapping frequency of the target tweezer or optical lattice site. What is left to show is how to implement $\mathrm{BRAID}$. We propose its realisation as a product $\mathrm{BRAID}_{ij}=\mathrm{PAIR^{\prime}}_{ij}\mathrm{\sqrt{iSWAP}}^{\mathrm{f}}_{ij}$, where $\mathrm{PAIR^{\prime}}_{ij}$ is the pairing gate
@@
To implement pairing, we propose the pulse-sequence in Fig. 5a. It employs photodissociation of homonuclear diatomic bosonic molecules consisting of the same atomic species as the fermions [58, 59], see Fig. 5b. Photodissociation realizes the boson-fermion gate $U_{\mathrm{diss}}=\exp\left(-i\frac{\pi}{4\sqrt{M}}\left(b_{i}p^{\dagger}_{i\uparrow}p^{\dagger}_{i\downarrow}+b^{\dagger}_{i}p_{i\downarrow}p_{i\uparrow}\right)\right)$, where $b$ is the bosonic annihilation operator of molecules (we assume there are $M$ of them) residing in the ground state of the tweezer and obeying commutation relations $[b_{i},b_{j}^{\dagger}]=\delta_{ij}$. $\uparrow$ and $\downarrow$ label two internal states of the atom, where so far we assumed that all atoms reside in $\downarrow$, i.e. $p_{i}=p_{i\downarrow}$. In order to convert this same-site gate into a non-local gate, we conjugate it with a fermion spin-selective tunneling $U_{\mathrm{t}\downarrow}(\theta)=\exp\left(-i\theta\left(p^{\dagger}_{i\uparrow}p_{j\uparrow}+h.c.\right)\right)$ and a fermionic spin-flip $U_{\uparrow\downarrow}(\theta)=\exp\left(-i\theta\left(p^{\dagger}_{i\downarrow}p_{i\uparrow}+h.c.\right)\right)$. To suppress interactions between the molecules and the fermions, the molecules reside in a separate tweezer when not in use for gates and are tunneled into the target tweezer by moving the two tweezers close to each other, implementing a beam-splitter operation $\mathrm{BS}(\theta)=\exp\left(\theta\left(b_{i}^{\dagger}b_{j}-h.c.\right)\right)$. Altogether, the sequence in Fig. 5a implements
$\mathrm{PAIR}_{ij}=\exp\left(i\frac{\pi}{4\sqrt{M}}\left(b_{k}p_{i}^{\dagger}p_{j}^{\dagger}+h.c.\right)\right).$ (9)
@@
We now discuss how we use Eq. (9) to effectively implement Eq. (8). If the molecules were in a coherent state with large amplitude $\sqrt{M}$, we could replace $b\rightarrow\braket{b}=\sqrt{M}$ and the two gates would be equal. However, in the case of atoms trapped in dipole traps, Fock states or mixtures thereof are more naturally prepared than coherent states. To overcome this conceptual challenge, we in the following use the fact that large-$M$ Fock states cannot be distinguished from coherent states by measurements [60].
@@
To show that we can indeed effectively implement Eq. (8) employing a large-$M$ Fock state of molecules, consider the thought experiment shown in Fig. 5c. The first pairing gate is supposed to create a “fermionic Bell-pair” (|$(\ket{00}+i\ket{11})/\sqrt{2}$. To probe whether there is coherence, a phase gate with phase $\theta$ is applied and a second $\mathrm{PAIR}$ gate rotates the measurement basis into the effective Bell basis. This realizes a Ramsey-type experiment. Simulating this experiment numerically, we find Ramsey fringes with a contrast that approaches unity as $M\rightarrow\infty$. This shows that, after the first pairing gate, there was indeed effective coherence between states |$\ket{00}$ and |$\ket{11}$. An intuitive way to understand this observation is that the molecule number distribution in each mode after the beamsplitter is approximately Poissonian—just like in a coherent state. The entanglement between both modes then leads to an effective phase coherence between the modes. This effective coherence is then revealed by the second pulse, which disentangles the fermions from the molecules. We say “effective” coherence because the reduced density matrix of the fermions is always diagonal; indeed, by measuring the total number of molecules, one projects the fermions onto |$\ket{00}$ or |$\ket{11}$. Therefore, it is important that the molecules are not measured while the fermions are entangled with them. In Fig. S1, we additionally perform state tomography numerically, showing that the state after the first pairing gate is indeed effectively a Bell pair.
@@
We probe the gate fidelity of our effective pairing gate in Fig. 5e using Choi state tomography, which we generalize to fermionic gates, see circuit shown in Fig. S1. We find that only $100$ molecules suffice to create a gate with $>99.8\,\%$ fidelity. In the Methods, we also show that even a mixed state with a Poissonian number distribution can be used instead of |$\ket{M}$. Non-trivially, Fig. 5c and Fig. 5d show that our scheme works even if the second pairing gate is not performed on the same molecule site, which enables full parallelizability of pairing operations. For that to be the case, the molecule tweezers need to be effectively coherent with respect to each other, just like lasers manipulating different qubits within a quantum computer need to be phase-locked. The cross-coherence of molecule tweezers is guaranteed by creating them from one common Fock state with a beamsplitter as done in Fig. 5c. To parallelize this further, many molecule tweezers can be created by repeatedly applying beamsplitters starting from one large Fock state, for example by tunneling from a single large dipole trap. While our proposal is challenging to implement, coherent dissociation of Feshbach molecules is routinely performed in cold-atom experiments [58].
@@
Our work opens the field of fermion-qubit error correction and fault-tolerant algorithms. Of prime importance is the error threshold of our codes for the strongly biased errors present in experiments. Moreover, fermionic versions of quantum low-density-parity check codes may be developed to enable a finite rate of logical fermions per codeblock. Due to our reliance on transversal gates, algorithmic fault tolerance can be explored within our approach [34].
@@
Our architecture might offer an advantageous approach not only for simulating fermions but also for quantum computations that are natively qubit-only. We encoded logical fermion operators in odd-weight fermion operators—instead, logical qubits can be encoded in even-weight logical operators, for example Majorana surface codes which have been shown to have high thresholds [72]. One advantage of this approach is that lost atoms are replenished by coupling to the molecule bath, offering an alternative to continuous atom reloading [55], which is one of the challenging aspects of neutral-atom quantum computing.
Finally, motivated by our depth improvement of the qubit algorithm for materials, development of fermion-hardware-inspired qubit algorithms might lead to further improvements in qubit algorithms for fermionic simulation.
@@ Acknowledgments
We acknowledge discussions with Ignacio Cirac, Daniel Gonzalez-Cuadra, Christopher Kang, Paul Julienne, Robert Ott, Shraddha Singh, Maryam Mudassar, Nathan Wiebe, Torsten Zache, Johannes Zeiher, and Hengyun Zhou.